ON THE NUMBER OF DAUBECHIES SCALING FUNCTIONS AND A CONJECTURE OF CHYZAK ET

AL.

Abstract. Using a result on Riesz factorizations, we show that there are at most 2^{N−1} and at least 2^{⌊N/2⌋} distinct Daubechies scaling functions with support in [1−N, N].